OpenStudy (anonymous):

Try to answer these brain teaser! Diophantus was a Greek mathematician who lived in the third century. He was one of the first mathematicians to use algebraic symbols. Most of what is known about Diophantus's life comes from an algebraic riddle from around the early sixth century. The riddle states: Diophantus's youth lasted one sixth of his life. He grew a beard after one twelfth more. After one seventh more of his life, he married. 5 years later, he and his wife had a son. The son lived exactly one half as long as his father, and Diophantus died four years after his son.

OpenStudy (anonymous):

The question is: How many years did Diophantus live?

OpenStudy (yttrium):

being one yr old is considered in his youth life, right?

OpenStudy (anonymous):

Umm, yeah...

OpenStudy (dumbcow):

are you ready.... 84

OpenStudy (yttrium):

\[\frac{ x }{ 6 }+\frac{ x }{ 12 }+\frac{ x }{ 7 }+5+\frac{ x }{ 2 }+4=x\]

OpenStudy (anonymous):

@dumbcow You got it right...!:)))

OpenStudy (yttrium):

using mental calculations, that will yield to 84

OpenStudy (yttrium):

more

OpenStudy (anonymous):

@Yttrium Try these one... You have a piece of paper, 10cm by 10cm. Area = 100cm^2. For some reason, you need a square piece of paper with an area of 50cm^2. Using the paper you have, what's an easy way of getting the new square?